let A,B,C,D,E be points on the complex plane representing the complex numbers z1,z2,z3,z4 and z5 respecctively. if  (z3 – z2) z4= (z1 –  z2) z5 .prove that the triangles ABC and DOE are similar.    (P.S – i think that it should be ∆EOD instead of ∆DOE , if not then please explain why )

Grade:Select Grade

1 Answers

arvind
7 years ago
The question is correct,since it is given  (z3 – z2) z4= (z1 –  z2) z5 ,it implies 
(z3- z2) / (z1-z2)=z5/z4 that is CB/BA=EO/DO or... CB/EO=BA/DO .
   Also since (z3- z2) / (z1-z2)=z5/z4 it implies argument[(z3- z2) / (z1-z2)] = argument[z5/z4], which means angle ABC = angle DOE
  Thus triangle ABC AND DOE ARE similar BY SAS TEST OF similarity.
HOPE THAT HELPED,anydoubt plz feel free to ask
Cheers!